OpenStudy (yodalicious):

PLEASE HELP!!! Write the equation of a hyperbola with vertices at (-4, 0) and (4, 0) and co-vertices (0, 5) and (0, -5).

OpenStudy (mathmale):

Hello. You may benefit from sketching this hyperbola. Through your sketch, you'll learn whether the hyperbola opens up and down (vertical) or to the right or left (horizontal). Next, you'll need to learn or review how the positions of the foci and vertices are related. I've done a quick Internet search and have come up with the following: https://www.google.com/search?sourceid=chrome-psyapi2&ion=1&espv=2&es_th=1&ie=UTF-8&q=equation%20of%20a%20hyperbola&oq=equation%20of%20a%20hyperbola&aqs=chrome..69i57j0l5.4975j0j7 By skimming through some of these search results, you'll find out what "a," "b" and "c" represent and how they are related to one another. Last hint (for now): The equation of a horiz. hyperbola with center at the origin (0,0) is \[\frac{ x^2 }{ a^2 }-\frac{ y^2 }{ b^2 }=1\]

OpenStudy (mathmale):

This may or many not be the actual equation you'll want to use.

OpenStudy (yodalicious):

thank you

OpenStudy (mathmale):

my pleasure!

OpenStudy (yodalicious):

oh okay